Calculate the standard cell potential at 25 ∘C for the reaction

X(s)+2Y+(aq)→X2+(aq)+2Y(s)

where ΔH∘ = -627 kJ and ΔS∘ = -153 J/K .
Answer #1

AG=-DF Eolel) Change AG free inerjy dectrors invalved Faraday 96500 (oloamls Fis oNe E ENF of the cellAG OHTAS AG=(-697 X 103 3)-(raK -1537 G -627000 +45594 -5 8। ५०6 ) Mo D = AGE -NEE e -581406 = - 500 X E C ऽ8।५०८ 2 X96500